# Equivariant CM minimization for extremal manifolds

### v1

@@ 1. Introduction
Let $C$ be a smooth projective curve with a special point 00, and let $(\mathcal{X},\mathcal{L})\to C$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\to C$ be fillings of a flat family $(X,L)\to C\setminus\{0\}$ such that the special fiber $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth, admits a cscK metric and has discrete automorphism group. Then
@@ 1.1. Main results
Let $R$ be a discrete valuation ring (DVR) with fraction field $K$, whose residue field is the field of complex numbers $\mathbb{C}$. The spectrum $\operatorname{Spec}R$ consists of a closed point 00 and a generic point $\operatorname{Spec}K$. The guiding example is the ring of formal power series $R=\mathbb{C}[\![t]\!]$, whose spectrum is called the formal disk and $K=\mathbb{C}(\!(t)\!)$ is the field of formal Laurent series.
Let $(X,L)$ be a polarized $K$-variety, which we think of as a degeneration of a polarized variety over $\operatorname{Spec}K$, and choose a maximal torus $T$ in the automorphism group $\mathrm{Aut}_{K}(X,L)$. The main result of the paper is the following generalization of Theorem 1.1, allowing automorphisms and extremal metrics (i.e. Kähler metrics $\omega$ satisfying ¯$\overline{\partial}\,\mathrm{grad}^{1,0}_{\omega}\mathrm{scal}(\omega)=0$, see [Cal82]):
@@ 1.2. Strategy of the proof
The second step is proving a new stability result involving the left-hand side of the Equation (1.1) in the case where $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth and extremal. More precisely, we prove that for all torus-invariant good filtrations $\mathcal{F}$, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$, where the inequality is strict whenever the reduced norm $||\mathcal{F}||_{T_{0}}$ is strictly positive. This generalizes works by Székelyhidi [Szé15], Hattori [Hat24] and Stoppa–Székelyhidi [SS11] for extremal and cscK manifolds with automorphisms.
@@ 2.1. The CM line bundle and models
$\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})=\frac{1}{2V(\mathcal{X}_{b},\mathcal{L}_{b})}\bigg{(}\lambda_{n}(\mathcal{X},\mathcal{L})-\frac{n\mu(\mathcal{X}_{b},\mathcal{L}_{b})}{n+1}\lambda_{n+1}(\mathcal{X},\mathcal{L})\bigg{)}$
@@ 2.1.1. Models over DVRs
We next discuss models, which will be the class of families of interest in the present paper. Let $R$ be a discrete valuation ring, endowed with a valuation $v$, with fraction field $K$ and residue field $\mathbb{C}$. Let $(X,L)$ be an $n$-dimensional normal $\mathbb{Q}$-polarized $K$-variety. This means that $(X,L)\to\operatorname{Spec}K$ is a flat family of $\mathbb{Q}$-polarized varieties of relative dimension $n$, where $L$ is a relatively ample $\mathbb{Q}$-line bundle.
@@
The CM line bundles of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are canonically isomorphic over the generic fiber, and using the identification with $(X,L)$, one may identify the difference of CM line bundles as a rational number, as follows: by functoriality of the Knudsen–Mumford expansion, the identification $(\mathcal{X}_{K},\mathcal{L}_{K})\simeq(X,L)\simeq(\mathcal{X}_{K}^{\prime},\mathcal{L}_{K}^{\prime})$ gives rise to a canonical isomorphism
$\big{(}\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\big{)}\big{|}_{\operatorname{Spec}K}\simeq\mathcal{O}_{\operatorname{Spec}K},$
of line bundles on $\operatorname{Spec}K$. This isomorphism may also be viewed as an isomorphism of $K$-vector spaces
$\varphi:H^{0}(\operatorname{Spec}K,\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))\to K,$
where the domain contains $\Lambda=H^{0}(\operatorname{Spec}R,\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})-\lambda_{\mathrm{CM}}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))\cong R$.
@@
The difference of CM degrees of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ of $(X,L)$ is defined as the valuation
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=v(\gamma),$
where $\gamma$ is any non-zero element $K$ such that $\gamma\cdot R=\varphi(\Lambda)$. We write $\mathrm{CM}(\mathcal{X},\mathcal{L})\geq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, whenever $\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\geq 0$, and similarly for $\leq$ and strict inequalities.
Within K-stability, the idea of associating a number to two line bundles over the formal disk endowed with a fixed identification away from the special fiber goes back to Donaldson [Don12] (see also Wang [Wan12]). In the case $R=\mathbb{C}[\![t]\!]$, the number $v(\gamma)$ coincides with the unique largest integer $\nu$ such that $t^{-\nu}\varphi(s|_{\operatorname{Spec}\mathbb{C}(\!(t)\!)})$ extends to a regular function on $\operatorname{Spec}\mathbb{C}[\![t]\!]$.
@@
compatible with the morphisms to $\operatorname{Spec}R$. Write $\mu^{*}\mathcal{L}=\mu^{\prime*}\mathcal{L}^{\prime}+E-a\mathcal{Y}_{0}$ for an effective $\mathbb{Q}$-divisor $E$ supported in $\mathcal{Y}_{0}$ and a rational number $a>0$. This gives rise to an embedding
@@
We may assume that $-E$ is $\mu$-relatively ample by replacing $(\mathcal{Y},\mathcal{M})$ with the ample model $(\operatorname{Proj}_{\mathcal{X}}(\bigoplus_{k\geq 0}\mu_{*}\mathcal{O}_{\mathcal{Y}}(k\mathcal{L})),\mathcal{O}(1))$ over $(\mathcal{X},\mathcal{L})$. There exists a closed subscheme $\mathcal{Z}$ in $\mathcal{X}$ supported in the special fiber such that the morphism $\mu:\mathcal{Y}\to\mathcal{X}$ is the blow-up of $\mathcal{X}$ along $\mathcal{Z}$ with exceptional divisor $rE$ for an integer $r\geq 1$. By [Laz17, Lemma 5.4.24], we obtain
@@
(2.2) $\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big{)}=b_{0}k^{n+1}+b_{1}k^{n}+O(k^{n-1})$
@@
We first claim that for all integers $i\geq 1$, the length of $H^{i}(\mathcal{Y},k\mathcal{M})$ as an $R$-module satisfies
(2.3) $\mathrm{length}_{R}H^{i}(\mathcal{Y},k\mathcal{M})=O(k^{n-1})$
as $k\to+\infty$. This was proven by Hattori [Hat24, Claim 1, Page 18] in the case where the family extends to a smooth projective curve. Indeed the same proof also applies for DVRs, as the Leray spectral sequence and Serre’s vanishing theorem hold over Noetherian rings.
We may assume that $R$ is regular by normalizing the base, pulling back the family and choosing an ample model; by functoriality of the Knudsen–Mumford expansion, the difference of CM degrees is unchanged, while Equation (2.2) is computed over the central fiber. In particular, one may choose a smooth proper curve $C$ and a point 00 in $C$, such that $R$ is isomorphic to the local ring $\mathcal{O}_{C,0}$ with $v=\operatorname{ord}_{0}$.
The identification of the models with $(X,L)$ over the generic fiber gives rise to an explicit trivialization of the vector bundle $E_{k}=\pi_{*}(k\mathcal{L})\otimes\pi^{\prime}_{*}(k\mathcal{L}^{\prime})^{\vee}$ away from zero, for sufficiently divisible $k$. Gluing this vector bundle on $\operatorname{Spec}K$ to the trivial bundle on the complement $C\setminus\{0\}$, we obtain vector bundle $\overline{E}_{k}$ on $C$ of rank $N_{k}=h^{0}(\mathcal{X}_{0},k\mathcal{L}_{0})$, satisfying
$\chi(C,\overline{E}_{k})=\chi_{R}(\operatorname{Spec}R,E_{k})+\chi(C,\mathcal{O}^{N_{k}}_{C})-1,$
where $\chi_{R}(-)=\mathrm{rank}_{R}H^{0}(-)+\sum_{i>0}(-1)^{i}\mathrm{length}_{R}H^{i}(-)$, by the inclusion-exclusion formula for Euler characteristics.
On one hand, by construction, the degree of $\overline{E}_{k}$ over $C$ coincides with the degree of $E_{k}$, defined as
$\deg E_{k}:=v(\det\gamma),$
where the trivialization away from zero gives rise to a canonical isomorphism $\varphi:H^{0}(\operatorname{Spec}K,E_{k})\to K^{N_{k}}$ and $\gamma$ is any element in $\operatorname{GL}_{N_{k}}(K)$ such that $\gamma\cdot R^{N_{k}}=\varphi(\Lambda_{k})$, for the $R$-lattice $\Lambda_{k}=H^{0}(\operatorname{Spec}R,E_{k})$. On the other hand, by Riemann-Roch for projective curves, for large $k$, we have
$\deg\overline{E}_{k}$ $=\chi(C,\overline{E}_{k})-\chi(C,\mathcal{O}^{N_{k}}_{C}),$
$=\chi_{R}(\operatorname{Spec}R,\pi_{*}(k\mathcal{L}))-\chi_{R}(\operatorname{Spec}R,\pi^{\prime}_{*}(k\mathcal{L}^{\prime}))-1,$
$=\chi_{R}(\mathcal{X},k\mathcal{L})-\chi_{R}(\mathcal{Y},k\mathcal{M}^{\prime})+O(k^{n-1}),$
$=\dim\!\big{(}H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big{)}+O(k^{n-1}),$
using Equation 2.3 and $\chi_{R}(\operatorname{Spec}R,\pi_{*}(k\mathcal{L}))=\chi_{R}(\mathcal{X},k\mathcal{L})$ for large $k$.
Applying the Knudsen–Mumford expansion to $E_{k}$ and taking the degree, we obtain
@@
$b_{1}$ $=-\frac{\deg(\lambda_{n}(\mathcal{X},\mathcal{L})-\lambda_{n}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))}{2n!},$
where the degree of the difference $\lambda_{i}(\mathcal{X},\mathcal{L})-\lambda_{i}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ is defined as for the difference of CM degrees. This yields the result. ∎
@@
With further hypotheses on the family $\mathcal{Y}\to\operatorname{Spec}R$, the difference of CM degrees may be defined in different ways. In the case where $\mathcal{Y}$ is normal and compactifies to a family over a smooth proper curve (still denoted $\mathcal{Y}$), we have
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=\frac{(\mu^{*}\mathcal{L}^{n}-\mathcal{M}^{n}).K_{\mathcal{Y}/R}-\frac{n\mu}{n+1}(\mu^{*}\mathcal{L}^{n+1}-\mathcal{M}^{n+1})}{2V}.$
If $\mathcal{Y}$ is smooth and the family $\mathcal{Y}\to\operatorname{Spec}R$ is the restriction of a family over a smooth (not necessarily proper) curve, the difference of CM degrees may also be viewed through Deligne pairings, as in this case the $\mathbb{Q}$-line bundles $\lambda_{i}$ in the Knudsen-Mumford expansion may be interpreted through Deligne pairings [PRS08]. Theorem 2.4 generalizes the formula for the Donaldson–Futaki invariant for test configurations and the difference of CM degrees for families over proper curves due to Odaka [Oda13] and Wang [Wan12].

### v2

@@ 1. Introduction
Let $C$ be a smooth projective curve with a special point 0, and let $(\mathcal{X},\mathcal{L})\to C$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\to C$ be fillings of a flat family $(X,L)\to C\setminus\{0\}$ such that the special fiber $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth, admits a cscK metric and has discrete automorphism group. Then
@@ 1.1. Main results
Let $R$ be a discrete valuation ring (DVR) with fraction field $K$, whose residue field is the field of complex numbers $\mathbb{C}$. The spectrum $\operatorname{Spec}R$ consists of a closed point 0 and a generic point $\operatorname{Spec}K$. The guiding example is the ring of formal power series $R=\mathbb{C}[\![t]\!]$, whose spectrum is called the formal disk and $K=\mathbb{C}(\!(t)\!)$ is the field of formal Laurent series.
Let $(X,L)$ be a polarized $K$-variety, which we think of as a degeneration of a polarized variety over $\operatorname{Spec}K$, and choose a maximal torus $T$ in the automorphism group $\mathrm{Aut}_{K}(X,L)$. The main result of the paper is the following generalization of Theorem 1.1, allowing automorphisms and extremal metrics (i.e. Kähler metrics $\omega$ satisfying $\overline{\partial}\,\mathrm{grad}^{1,0}_{\omega}\mathrm{scal}(\omega)=0$, see [Cal82]):
@@ 1.2. Strategy of the proof
The second step is proving a new stability result involving the left-hand side of the Equation (1.1) in the case where $(\mathcal{X}_{0},\mathcal{L}_{0})$ is smooth and extremal. More precisely, we prove that for all torus-invariant good filtrations $\mathcal{F}$, we have $\mathrm{DF}_{T_{0}}(\mathcal{F})\geq 0$, where the inequality is strict whenever the reduced norm $\mathcal{F}||_{T_{0}}$ is strictly positive. This generalizes works by Székelyhidi [Szé15], Hattori [Hat24] and Stoppa–Székelyhidi [SS11] for extremal and cscK manifolds with automorphisms.
@@ 2.1. The CM line bundle and models
$\lambda_{\mathrm{CM}}(\mathcal{X},\mathcal{L})=\frac{1}{2V(\mathcal{X}_{b},\mathcal{L}_{b})}\bigg(\lambda_{n}(\mathcal{X},\mathcal{L})-\frac{n\mu(\mathcal{X}_{b},\mathcal{L}_{b})}{n+1}\lambda_{n+1}(\mathcal{X},\mathcal{L})\bigg),$
@@ 2.1.1. Models over DVRs
We next discuss models, which will be the class of families of interest in the present paper. Let $R$ be a discrete valuation ring with fraction field $K$ and residue field $\mathbb{C}$, and choose a valuation $\mathsf{v}:K^{*}\to\mathbb{R}$. Let $(X,L)$ be an $n$-dimensional normal $\mathbb{Q}$-polarized $K$-variety. This means that $X\to\operatorname{Spec}K$ is a flat morphism of relative dimension $n$, and $L$ is a relatively ample $\mathbb{Q}$-line bundle on $X$.
@@
The CM line bundles $\lambda_{\mathrm{CM}}$ and $\lambda^{\prime}_{\mathrm{CM}}$ of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ are canonically isomorphic over the generic fiber, and using the identification with $(X,L)$, one may identify the difference of CM line bundles as a rational number, as follows: by functoriality of the Knudsen–Mumford expansion, the identification $(\mathcal{X}_{K},\mathcal{L}_{K})\simeq(X,L)\simeq(\mathcal{X}_{K}^{\prime},\mathcal{L}_{K}^{\prime})$ gives rise to an explicit isomorphism
$\Phi:\lambda_{\mathrm{CM}}|_{\operatorname{Spec}K}\to\lambda_{\mathrm{CM}}^{\prime}|_{\operatorname{Spec}K}$
of line bundles. This isomorphism may also be viewed as an isomorphism of $K$-vector spaces
$\Phi_{*}:H^{0}(\operatorname{Spec}K,m\lambda_{\mathrm{CM}})\to H^{0}(\operatorname{Spec}K,m\lambda_{\mathrm{CM}}^{\prime}),$
where $m$ is any sufficiently divisible integer such that one takes sections of line bundles. The domain and target contain $\Lambda=H^{0}(\operatorname{Spec}R,m\lambda_{\mathrm{CM}})$ and $\Lambda^{\prime}=H^{0}(\operatorname{Spec}R,m\lambda_{\mathrm{CM}}^{\prime})$ as $R$-submodules, respectively.
@@
The difference of CM degrees of two models $(\mathcal{X},\mathcal{L})$ and $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$ of $(X,L)$ is the valuation
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}):=m^{-1}\mathsf{v}(\gamma),$
where $\gamma$ is any non-zero element $K$ such that $\Phi_{*}(\Lambda)=\gamma\cdot\Lambda^{\prime}$. We write $\mathrm{CM}(\mathcal{X},\mathcal{L})\geq\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, whenever $\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})\geq 0$, and similarly for other inequalities.
Within K-stability, the idea of associating a number to two line bundles over the formal disk endowed with a fixed identification away from the special fiber goes back to Donaldson [Don12] (see also Wang [Wan12]). In the case $R=\mathbb{C}[\![t]\!]$, the number $\mathsf{v}(\gamma)$ coincides with the unique largest integer $\nu$ such that, giving a trivializing section $s$ of $\Lambda$, the section $t^{-\nu}\Phi_{*}s$ extends to a section on $\operatorname{Spec}\mathbb{C}[\![t]\!]$. In this case, the difference of CM degrees is equal to the degree of the line bundle $\overline{\lambda_{\mathrm{CM}}-\lambda_{\mathrm{CM}}^{\prime}}$ on $\mathbb{P}^{1}$ obtained by gluing the line bundle $\lambda_{\mathrm{CM}}-\lambda_{\mathrm{CM}}^{\prime}$ on $\operatorname{Spec}\mathbb{C}[\![t]\!]$ to the trivial line bundle on $\mathbb{P}^{1}\setminus 0$, using the explicit identification $\lambda_{\mathrm{CM}}-\lambda_{\mathrm{CM}}^{\prime}|_{\operatorname{Spec}\mathbb{C}(\!(t)\!)}\simeq\mathcal{O}_{\operatorname{Spec}\mathbb{C}(\!(t)\!)}$ on the intersection.
@@
compatible with the morphisms to $\operatorname{Spec}R$. After potentially changing $(\mathcal{X},\mathcal{L})$ with $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$, we can write $\mu^{*}\mathcal{L}=\mu^{\prime*}\mathcal{L}^{\prime}+E-a\mathcal{Y}_{0}$ for an effective $\mathbb{Q}$-divisor $E$ supported in $\mathcal{Y}_{0}$ and a rational number $a>0$. This gives rise to an embedding
@@
Assume that E-E is $\mu$-relatively ample by replacing $(\mathcal{Y},\mathcal{M})$ with the ample model $(\operatorname{Proj}_{\mathcal{X}}(\bigoplus_{k\geq 0}\mu_{*}\mathcal{O}_{\mathcal{Y}}(k\mathcal{L})),\mathcal{O}(1))$ over $(\mathcal{X},\mathcal{L})$. There exists a closed subscheme $\mathcal{Z}$ in $\mathcal{X}$ supported in the special fiber such that the morphism $\mu:\mathcal{Y}\to\mathcal{X}$ is the blow-up of $\mathcal{X}$ along $\mathcal{Z}$ with exceptional divisor $rE$ for an integer $r\geq 1$. By [Laz17, Lemma 5.4.24], we obtain
@@
(2.2) $\dim_{\mathbb{C}}\big(H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big)=b_{0}k^{n+1}+b_{1}k^{n}+O(k^{n-1})$
@@
We may assume that $R$ is regular by normalizing the base, pulling back the family and choosing an ample model; as the Knudsen–Mumford expansion is functorial, the difference of CM degrees is unchanged, while Equation (2.2) is computed over the central fiber.
Write $E_{k}=\pi_{*}(k\mathcal{L})$ as a vector bundle on $\operatorname{Spec}R$, where $\pi:\mathcal{X}\to\operatorname{Spec}R$ is the structure morphism, as write
$H_{k}^{i}$ $=H^{i}(\operatorname{Spec}K,E_{k}),$
$H_{k,R}^{i}$ $=H^{i}(\operatorname{Spec}R,E_{k}).$
Similarly, write $E_{k}^{\prime}$, $H_{k}^{\prime\,i}$ and $H_{k,R}^{\prime\,i}$ for the same objects, but using the model $(\mathcal{X}^{\prime},\mathcal{L}^{\prime})$. Similar to Definition 2.3, one can define the degree of the vector bundle $E_{k}\otimes E_{k}^{\prime\,\vee}$ with respect to the identification $E_{k}|_{\operatorname{Spec}K}\simeq E^{\prime}_{k}|_{\operatorname{Spec}K}$, explicitly given by the models, as the valuation
$\deg\big(E_{k}\otimes E_{k}^{\prime\,\vee}\big):=\mathsf{v}(\det\gamma),$
where $\gamma$ is any element in $\operatorname{GL}_{K}(H_{k}^{\prime\,0})$ such that $\Phi(H_{k,R}^{0})=\gamma(H_{k,R}^{\prime\,0})$, where $\Phi:H_{k}^{i}\to H_{k}^{\prime\,i}$ is the isomorphism associated to the identification. Note that by construction, we have
(2.3) $\deg(\det E_{k}-\det E_{k}^{\prime})=\deg(E_{k}\otimes E_{k}^{\prime\,\vee}),$
where the degree of the difference of the determinant line bundles is defined the same way as for the difference of CM line bundles.
The localized Euler characteristic of $E_{k}\otimes E_{k}^{\prime\,\vee}$ with respect to the identification is defined as the difference of lengths or $R$-modules
$\chi\big(E_{k}\otimes E_{k}^{\prime\,\vee}\big)$ $=\sum_{i\geq 0}(-1)^{i}\,\mathrm{len}_{R}\big(H_{k}^{i}/(H_{k,R}^{i}\cap\Phi^{-1}(H_{k,R}^{\prime\,i})\big).$
Note that $H^{i}_{k}=0$ for all $i>0$ and $k$ very large, as $\operatorname{Spec}R$ is affine, and that $\Phi^{-1}(H_{k,R}^{\prime\,0})=H^{0}(\operatorname{Spec}R,\pi_{*}^{\mathcal{Y}}(k\mathcal{M}))=:H^{\mathcal{Y}}_{k,R}$, by construction of $\mathcal{Y}$, where $\pi^{\mathcal{Y}}:\mathcal{Y}\to\operatorname{Spec}R$ is the structure morphism.
By the localized Riemann–Roch formula over $\operatorname{Spec}R$ (see for example [Ful13, Theorem 18.2], where we apply the theory to the closed subscheme 0 in $\operatorname{Spec}R$), we then have
$\deg\big(E_{k}\otimes E_{k}^{\prime\,\vee}\big)$ $=\chi\big(\pi_{*}(k\mathcal{L})\otimes\pi^{\prime}_{*}(k\mathcal{L}^{\prime})^{\vee}\big),$
$=\mathrm{len}_{R}\big(H_{k}^{0}/H^{\mathcal{Y}}_{k,R}\big),$
$=\dim_{\mathbb{C}}\big(H_{k,R}^{0}/H^{\mathcal{Y}}_{k,R}\big),$
$=\dim_{\mathbb{C}}\big(H^{0}(\mathcal{X},k\mathcal{L})/H^{0}(\mathcal{Y},k\mathcal{M})\big),$
for $k$ very large, where we used [Har13, Proposition 8.5] in the last line. Applying the Knudsen–Mumford expansion to Equation (2.3) and taking the degree, we obtain
@@
$b_{1}$ $=-\frac{\deg(\lambda_{n}(\mathcal{X},\mathcal{L})-\lambda_{n}(\mathcal{X}^{\prime},\mathcal{L}^{\prime}))}{2n!}.$
This yields the result. ∎
@@
With further hypotheses on the family $\mathcal{Y}\to\operatorname{Spec}R$, the difference of CM degrees may be defined in different ways. In the case where the models are normal and compactify to a family over a smooth proper curve $C$, we have
$\mathrm{CM}(\mathcal{X},\mathcal{L})-\mathrm{CM}(\mathcal{X}^{\prime},\mathcal{L}^{\prime})=\frac{(\mu^{*}\mathcal{L}^{n}-\mathcal{L}^{\prime\,n}).K_{\mathcal{X}/C}-\frac{n\mu}{n+1}(\mu^{*}\mathcal{L}^{n+1}-\mathcal{L}^{\prime\,n+1})}{2V}.$
If the models are smooth and the family are restrictions of famililies over a smooth (not necessarily proper) curve, the difference of CM degrees may also be viewed through Deligne pairings, as in this case the $\mathbb{Q}$-line bundles $\lambda_{i}$ in the Knudsen-Mumford expansion may be interpreted through Deligne pairings [PRS08]. Theorem 2.4 generalizes the formula for the Donaldson–Futaki invariant for test configurations and the difference of CM degrees for families over proper curves due to Odaka [Oda13] and Wang [Wan12].